1.How did you decide how to design your spinner? In other words, how did you go about planning the design?
40% of the spinner needed to be red, 40% needed to be white, and 20% needed to be blue. In order for 100% of the spins to land on red or odd (P(red or odd)) the white and blue portion of the spinner would have to be odd numbers. The next step is to make it possible for the spinner to land on blue and 3 (P(blue and 3)) by putting the number 3 on the 20% section that is blue.
2.How did you know that your spinner design met the criteria above? Show evidence.
If 20% of the spinner is the number 3, 40% is 9, and 40% is red, the chance of landing on red or an odd number is 100% (20%+40%+40%=100%.) Since the blue portion is 20%, if the number 3 is on this portion 20% of the spinner is blue and 3.
The results prove that about 20% of the time the spinner lands on blue (3) and 100% of the time the spinner will land on either red, blue (odd number), or white (odd number.)


No comments:
Post a Comment